OpenStudy (anonymous):

If f(x) = −2x + 1, find f(−6). 12 13 −11 −7

myininaya (myininaya):

Where there is x, put -6

myininaya (myininaya):

-2(-6)+1 Use order of operations to simplify

OpenStudy (ledah):

Here's an example This is simple substitution. Where you see an x put in the number 3 f(3) = 2(3) - 1 f(3) = 6 - 1 f(3) = 5 <----------Answer

myininaya (myininaya):

For this problem that @Ledah has, she (is that right? ) taking f to be f(x)=2x-1

myininaya (myininaya):

So for f(x)=2x-1 we can do another example What is f(-5) So we are going to replace x with -5 2(-5)-1 Using order of operations we do multiplication first -10-1 Now -10-1=-11 So f(-5)=-11

OpenStudy (ledah):

hehe I'm a male :P